Express 1 / y(y+1) in partial fractions.

Hence solve the differential equation dy/dx = y(y+1) / x given that y=4 when x=2 expressing y explicitly in terms of x.


1
Expert's answer
2021-09-21T00:21:23-0400

Solution:

"\\ \dfrac1{y(y+1)}=\dfrac{A}{y}+\dfrac{B}{y+1}
\\\Rightarrow \dfrac1{y(y+1)}=\dfrac{A(y+1)+By}{y(y+1)}
\\\Rightarrow 1=Ay+A+By
\\\Rightarrow 0.y+1=y(A+B)=A"

On comparing both sides,

"A+B=0,A=1
\\\Rightarrow B=-1"

Thus, "\dfrac1{y(y+1)}=\dfrac{1}{y}-\dfrac{1}{y+1}" ...(i)

Now, given "\dfrac{dy}{dx}=\dfrac{y(y+1)}{x}"

"\Rightarrow \dfrac{1}{y(y+1)}dy=\dfrac{1}{x}dx
\\ \Rightarrow [\dfrac{1}{y}-\dfrac{1}{y+1}]dy=\dfrac{1}{x}dx \ [using\ (i)]"

On integrating both sides,

"\ln |y|-\ln|y+1|=\ln |x|+\ln C
\\\Rightarrow \ln|\dfrac{y}{y+1}|=\ln |\dfrac xC|
\\\Rightarrow \dfrac{y}{y+1}=\dfrac xC"

Now, put y=4, x=2

"\dfrac{4}{5}=\dfrac 2C
\\\Rightarrow C=\dfrac52"

Thus, the solution is "\dfrac{y}{y+1}=\dfrac {2x}{5}"